find the functions f and g so that f of g =H. H(x)=(6x+9)^4
please show work so i understand this better. thank you

Answer by stanbon(75887)

find the functions f and g so that f of g =H. H(x)=(6x+9)^4
-------
H(x) = fog(x)
Notice that g is the 1st function to work on "x".
Let g(x) = 6x+9
----
Notice that f does its work on g(x)
Let f(x) = x^4
----
Then H(x) = fog(x) = f[6x+9] = (6x+9)^4
==============================================
Cheers,
Stan H.